### v1

@@ 4.2.2 PEMC for HJM
with $\sigma_{0}$ and $\alpha_{\sigma}$ as part of $\bm{\theta}$. Similarly, for a baseline initial forward curve, we use:
@@
with $f_{0},c_{f}$, and $\alpha_{f}$ part of the $\bm{\theta}$. While this analytical form serves as our data-generating process, PEMC accesses it only through its discrete grid values with added noise. This approach mirrors our treatment of the local volatility surface in the previous section, where we used the parametric form in Carmona (2007) solely as a realistic baseline for generating training data.
Indeed, while HJM and these analytical forms (17)-(18) are formulated in continuous time, in practice we need to implement numerical discretization. Following the scheme in Glasserman (2013b), one discretizes both the time axis and the set of maturities i.e., time steps $\mathcal{T}_{t}=\{t_{0},t_{1},\ldots,t_{N_{T}}\}$ and maturity points $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{N_{M}}\}$ forming a time-maturity grid. In our experiment, for simplicity, we assume they share one grid $\mathcal{T}$ of size $|\mathcal{T}|$. Then, when sampling $\bm{\theta}\sim\bm{\Theta}$, one first sample $\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}$, then one sample the grids $\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}$ and $\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}$ from (17) and (18) with added noise $\mathcal{N}(0,(\frac{\sigma_{0}}{2(t_{j}+5)})^{2})$ and $\mathcal{N}(0,(\frac{1}{100(t_{j}+5)})^{2})$ respectively, on top of their baseline values. The noise level $\xi=\frac{\sigma_{0}}{2(t+5)}$. Paths are then generated using the discretization scheme described in Glasserman (2013b) (e.g., ${f}(t_{i},t_{j})={f}(t_{i-1},t_{j})\;+\;{\mu}(t_{i-1},t_{j})(t_{i}-t_{i-1})\;+\;\sqrt{t_{i}-t_{i-1}}{\sigma}(t_{i-1},t_{j})\mathcal{N}(0,1)$ where ${\mu}(t_{i-1},t_{j})$ is the discretized drift term determined through (16)). We refer interested readers to Glasserman (2013b) for the complete derivation and implementation details of this standard simulation scheme.
For PEMC, our the parameter $\bm{\theta}\sim\bm{\Theta}$ includes the volatility parameters $(\sigma_{0},\alpha_{\sigma})$ in (17) and initial forward curve parameters $(f_{0},c,\alpha_{f})$ in (18). The swaption’s fixed rate is set as $R=\exp(-\frac{\sum_{i}^{n-1}f(0,t^{\prime}_{i})}{T_{\text{final}}-t^{\prime}_{0}})$, reflecting the common practice where swap rates are typically determined in reference to the prevailing forward rate curve rather than being arbitrarily chosen. Here the parameters are notional amount $C$, start time $t^{\prime}_{0}$, payment interval $\Delta t^{\prime}$, and number of payments $n_{p}$. The simulation uses time step $\Delta t$ up to final maturity $T_{\text{final}}$. During training, parameters are sampled uniformly from the ranges specified in Table 9 where the evaluation $\bm{\theta}$ is also listed.
Table 9: Parameter Setup in HJM Model
@@
$\text{feature}=\{\underbrace{\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}}_{\text{volatility structure}},\underbrace{\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}}_{\bm{\theta}_{\text{model}}},\underbrace{\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}}_{{\text{initial forward curve}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}}\}.$
To handle both the two-dimensional grid of volatility structure $\sigma(t,T)$ and the one dimensional grid of initial forward curve $f(0,T)$, we design a three-branch neural network architecture, illustrated in Figure 2. The first branch, labeled“2D Function Encoder”, processes the 2d volatility structure grid using a CNN architecture with two 2D convolutional layers, each followed by batch normalization. The branch concludes with an average pooling operation and produces an embedding. The second branch processes the initial forward curve $f(0,T)$ grid through a “1D Function Encoder” utilizing 1D convolutional layers - a natural choice for sequential data (LeCun et al. 1998) - followed by batch normalization and average pooling to produce another embedding. The third branch handles the remaining input through fully connected layers with batch normalization. Finally, these three separate embeddings are then fed into a “Synthesizer” module that combines the information through multiple fully connected layers with batch normalization, ultimately producing the final prediction. This architecture effectively leverages both the spatial structure of the volatility surface through 2D CNNs (Simonyan and Zisserman 2014), the sequential nature of the forward curve through 1D CNNs (Oord et al. 2016), and the scalar parameters through standard deep learning techniques (He et al. 2016). The complete network architecture is detailed in Table 10.
Figure 2: Neural network architecture for modeling Swaption payoff.
@@
Table 10: Hyper-parameter setup for the neural network
Finally, following the same evaluation methodology, we assess PEMC’s performance in the HJM swaption pricing setting across sample sizes $n\in\{1000,3000,5000,7000,9000,11000\}$ with $N=10n$, using $5\times 10^{7}$ MC samples to establish the ground truth value. As presented in Table 11, the outcomes from 200 independent experiments confirm that PEMC remains highly effective within the context of interest rate derivatives. Despite the added complexity of managing both volatility structures and forward rate curves, PEMC achieves a 45-50% reduction in RMSE to standard Monte Carlo methods. This performance aligns with the variance reduction levels observed in our earlier examples. The boxplot is shown in Figure 4.2.2. The consistent effectiveness of PEMC across various financial instruments and modeling frameworks further underscores its versatility as a robust variance reduction technique.
Table 11: Root Mean Squared Error from 200 Experiments
@@
Boxplots for HJM Experiments.
4.3 Discussions and Extensions
@@
We conclude with several practical observations and potential extensions of the PEMC framework.
4.3.1 Evaluation Metric
@@
First, while training the neural network estimator $g$ using mean squared error (MSE) loss is common, it is not always clear how to interpret the resulting MSE score. Unlike some well-established benchmarks (e.g., classification accuracy), there is no canonical threshold or known “good” MSE value for a given problem. This ambiguity makes it challenging to determine when the network is sufficiently trained. To address this, we can exploit the fact that $g$ is meant to represent the conditional expectation $g=\mathbb{E}[f\mid\text{input}]$. If the network approximates this expectation well, then the sample average of $g(\bm{X})$ should be close to the sample average of $f(\bm{Y})$ over a given dataset. One practical diagnostic is to compute the Mean Absolute Relative Error (MARE) between these two averages. If $\mathbb{E}[g(\bm{X})]\approx\mathbb{E}[f(\bm{Y})]$, it provides a tangible indication that $g$ is capturing the underlying expectation. Our empirical experience suggests this criterion is very effective in practice (a 5-1% MARE typically indicates exceptional PEMC), complementing common techniques like early stopping in machine learning workflows. More importantly, as the we have shown in the theory, if $g$ closely approximates the conditional expectation, the variance reduction in PEMC is guaranteed—even if marginally—relative to standard Monte Carlo.
4.3.2 XVA, Greeks and Quasi-Monte Carlo
@@
Beyond pricing exotic options, this approach naturally extends to other computationally intensive Monte Carlo settings in quantitative finance. Notably, adjustments like Credit valuation adjustment (CVA) and more complex XVAs often require vast simulation runs, sometimes taking days to complete (Green et al. 2015, Andersen et al. 2017). Integrating PEMC with a well-trained neural network can cut these computational times drastically while preserving accuracy.
@@
Moreover, in the pricing of exotic derivatives and other path-dependent instruments, automatic differentiation (AD) techniques are often used to compute Greeks efficiently. By leveraging modern deep learning frameworks—such as PyTorch or TensorFlow—that support automatic differentiation natively, we can easily differentiate the trained network $g$ with respect to input parameters (Giles and Glasserman 2006). This enables quick and accurate sensitivity analysis without resorting to finite-difference approximations or nested simulations, which could lead to another direction of extension for PEMC.

### v2

@@ 5.2.1. PEMC for HJM
with $\sigma_{0}$ and $\alpha_{\sigma}$ as part of $\boldsymbol{\theta}$. Similarly, for a baseline initial forward curve, we use:
@@
with $f_{0},c_{f}$, and $\alpha_{f}$ part of the $\boldsymbol{\theta}$. While this analytical form serves as our data-generating process, PEMC accesses it only through its discrete grid values with added noise. This approach mirrors our treatment of the local volatility surface in the previous section, where we used the parametric form in [34] solely as a realistic baseline for generating training data.
Indeed, while HJM and these analytical forms (16)-(17) are formulated in continuous time, in practice we need to implement numerical discretization. Following the scheme in [60], one discretizes both the time axis and the set of maturities i.e., time steps $\mathcal{T}_{t}=\{t_{0},t_{1},\ldots,t_{N_{T}}\}$ and maturity points $\mathcal{T}=\{t_{1},t_{2},\ldots,t_{N_{M}}\}$ forming a time-maturity grid. In our experiment, for simplicity, we assume they share one grid $\mathcal{T}$ of size $|\mathcal{T}|$. Then, when sampling $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$, one first sample $\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}$, then one sample the grids $\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}$ and $\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}$ from (16) and (17) with added noise $\mathcal{N}(0,(\frac{\sigma_{0}}{2(t_{j}+5)})^{2})$ and $\mathcal{N}(0,(\frac{1}{100(t_{j}+5)})^{2})$ respectively, on top of their baseline values. The noise level $\xi=\frac{\sigma_{0}}{2(t+5)}$. Paths are then generated using the discretization scheme described in [60] (e.g., ${f}(t_{i},t_{j})={f}(t_{i-1},t_{j})\;+\;{\mu}(t_{i-1},t_{j})(t_{i}-t_{i-1})\;+\;\sqrt{t_{i}-t_{i-1}}{\sigma}(t_{i-1},t_{j})\mathcal{N}(0,1)$ where ${\mu}(t_{i-1},t_{j})$ is the discretized drift term determined through (15)). We refer interested readers to [60] for the complete derivation and implementation details of this standard simulation scheme.
For PEMC, our the parameter $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$ includes the volatility parameters $(\sigma_{0},\alpha_{\sigma})$ in (16) and initial forward curve parameters $(f_{0},c,\alpha_{f})$ in (17). The swaption’s fixed rate is set as $R=\exp(-\frac{\sum_{i}^{n-1}f(0,t^{\prime}_{i})}{T_{\text{final}}-t^{\prime}_{0}})$, reflecting the common practice where swap rates are typically determined in reference to the prevailing forward rate curve rather than being arbitrarily chosen. Here the parameters are notional amount $C$, start time $t^{\prime}_{0}$, payment interval $\Delta t^{\prime}$, and number of payments $n_{p}$. The simulation uses time step $\Delta t$ up to final maturity $T_{\text{final}}$. During training, parameters are sampled uniformly from the ranges specified in Table 4 where the evaluation $\boldsymbol{\theta}$ is also listed.
Table 4. Parameter Setup in HJM Model
@@
$\text{feature}=\{\underbrace{\{\sigma(t_{i},t_{j})\}_{t_{i}\leq t_{j},t_{i},t_{j}\in\mathcal{T}}}_{\text{volatility structure}},\underbrace{\sigma_{0},\alpha_{\sigma},f_{0},c_{f},\alpha_{f}}_{\boldsymbol{\theta}_{\text{model}}},\underbrace{\{f(0,t_{i})\}_{t_{i}\in\mathcal{T}}}_{{\text{initial forward curve}}},\underbrace{(W^{S}_{T},W^{\nu}_{T})}_{\mathbf{X}}\}.$
To handle both the two-dimensional grid of volatility structure $\sigma(t,T)$ and the one dimensional grid of initial forward curve $f(0,T)$, we design a three-branch neural network architecture, illustrated in Figure 3. The first branch, labeled“2D Function Encoder", processes the 2d volatility structure grid using a CNN architecture with two 2D convolutional layers, each followed by batch normalization. The branch concludes with an average pooling operation and produces an embedding. The second branch processes the initial forward curve $f(0,T)$ grid through a “1D Function Encoder" utilizing 1D convolutional layers - a natural choice for sequential data [83] - followed by batch normalization and average pooling to produce another embedding. The third branch handles the remaining input through fully connected layers with batch normalization. Finally, these three separate embeddings are then fed into a “Synthesizer" module that combines the information through multiple fully connected layers with batch normalization, ultimately producing the final prediction. This architecture effectively leverages both the spatial structure of the volatility surface through 2D CNNs [118], the sequential nature of the forward curve through 1D CNNs [101], and the scalar parameters through standard deep learning techniques [64]. The complete network architecture is detailed in Table 5.
Figure 3. Neural network architecture for modeling Swaption payoff.
@@
Table 5. Hyper-parameter setup for the neural network
Finally, following the same evaluation methodology, we assess PEMC’s performance in the HJM swaption pricing setting across sample sizes $n\in\{1000,3000,5000,7000,9000,11000\}$ with $N=10n$, using $5\times 10^{7}$ MC samples to establish the ground truth value. As presented in Table 6, the outcomes from 200 independent experiments confirm that PEMC remains highly effective within the context of interest rate derivatives. Despite the added complexity of managing both volatility structures and forward rate curves, PEMC achieves a 45-50% reduction in RMSE to standard Monte Carlo methods. This performance aligns with the variance reduction levels observed in our earlier examples. The boxplot is shown in Figure 4. The consistent effectiveness of PEMC across various financial instruments and modeling frameworks further underscores its versatility as a robust variance reduction technique.
Table 6. Root Mean Squared Error from 200 Experiments
@@
Figure 4. Boxplots for HJM Experiments.
5.3. Discussions, Extensions and More Examples
@@
We conclude with several observations and potential extensions of the PEMC framework.
5.3.1. Evaluation Metric
@@
First, while training the neural network estimator $g$ using mean squared error (MSE) loss is common, it is not always clear how to interpret the resulting MSE score. Unlike some well-established benchmarks (e.g., classification accuracy), there is no canonical threshold or known “good” MSE value for a given problem. This ambiguity makes it challenging to determine when the network is sufficiently trained. To address this, we can exploit the fact that $g$ is meant to represent the conditional expectation $g=\mathbb{E}[f\mid\text{input}]$. If the network approximates this expectation well, then the sample average of $g(\boldsymbol{X})$ should be close to the sample average of $f(\boldsymbol{Y})$ over a given dataset. One practical diagnostic is to compute the Mean Absolute Relative Error (MARE) between these two averages. If $\mathbb{E}[g(\boldsymbol{X})]\approx\mathbb{E}[f(\boldsymbol{Y})]$, it provides a tangible indication that $g$ is capturing the underlying expectation. Our empirical experience suggests this criterion is very effective in practice (a 5-1% MARE typically indicates exceptional PEMC), complementing common techniques like early stopping in machine learning workflows. More importantly, as the we have shown in the theory, if $g$ closely approximates the conditional expectation, the variance reduction in PEMC is guaranteed—even if marginally—relative to standard Monte Carlo.
5.3.2. XVA, Greeks and Quasi-Monte Carlo
@@
Beyond pricing exotic options, this approach naturally extends to other computationally intensive Monte Carlo settings in quantitative finance. Notably, adjustments like Credit valuation adjustment (CVA) and more complex XVAs often require vast simulation runs, sometimes taking days to complete [62, 4]. Integrating PEMC with a well-trained neural network can cut these computational times drastically while preserving accuracy.
@@
Moreover, in the pricing of exotic derivatives and other path-dependent instruments, automatic differentiation (AD) techniques are often used to compute Greeks efficiently. By leveraging modern deep learning frameworks—such as PyTorch or TensorFlow—that support automatic differentiation natively, we can easily differentiate the trained network $g$ with respect to input parameters [59]. This enables quick and accurate sensitivity analysis without resorting to lattice-structure approximations [133, 132] or nested simulations, which could lead to another direction of extension for PEMC.